OpenStudy (baby456):

Assuming x and y are positive,use properties of logarithms to write the expression as a sum or difference of a logarithm or multiples of logarithms. In cubed(x)/cubed(y)?

OpenStudy (baby456):

@jim_thompson5910

OpenStudy (baby456):

@zepdrix

OpenStudy (baby456):

@mathmale

jhonyy9 (jhonyy9):

do you think it hence ln x^3 /y^3 = ?

OpenStudy (baby456):

cubed root

jhonyy9 (jhonyy9):

do you know this property of logarithms log a -log b = log (a/b) log a^x = xlog a log a +log b = log a*b

OpenStudy (baby456):

yep

jhonyy9 (jhonyy9):

ok. so than just use the necessary

jhonyy9 (jhonyy9):

do you mean cuberoot of x divide cuberoot y ?

OpenStudy (baby456):

so the quoitent rule

OpenStudy (baby456):

yes

jhonyy9 (jhonyy9):

\[\ln \sqrt[3]x / \sqrt[3]{y} =\]

jhonyy9 (jhonyy9):

in this way ?

OpenStudy (baby456):

inx/3-iny/3

jhonyy9 (jhonyy9):

hence ? just ln x/3 -ln y/3 = ?

jhonyy9 (jhonyy9):

cubed x dont mean x^3 ?

jhonyy9 (jhonyy9):

just bc. you wrote in the text of your exercise cubed x and cubed y

OpenStudy (baby456):

ok so is it 1/3(in x- In y)

jhonyy9 (jhonyy9):

this is just than you wan saying that your exercise is ln x^(1/3) -ln y^(1/3) what is equal ln (cuberoot x / cuberoot y )

OpenStudy (baby456):

3

jhonyy9 (jhonyy9):

what mean this 3 ?

OpenStudy (baby456):

i dont know what your are asking

jhonyy9 (jhonyy9):

please clarify how is this your exercise correct ?

jhonyy9 (jhonyy9):

please make a copy past

OpenStudy (baby456):

when you use the quotient rule your trying to subtract

OpenStudy (baby456):

so ration cube root (x)= 1/3

OpenStudy (baby456):

so 1/3(In x- In y)

jhonyy9 (jhonyy9):

do you think ln a^x = xln a ln a^x -ln b^x = xln (a/b)

jhonyy9 (jhonyy9):

ok. ?

jhonyy9 (jhonyy9):

cuberoot x = x^(1/3)

jhonyy9 (jhonyy9):

do you think it so ?

OpenStudy (baby456):

yes

OpenStudy (baby456):

thats what my teacher taught us is the wayy i am solving it

jhonyy9 (jhonyy9):

so in this case there are ln (cuberoot x / cuberoot y ) = ln [x^(1/3) / y^(1/3) ] = ln ( x/y)^(1/3) = (1/3)ln x/y do you think it hence ?

OpenStudy (baby456):

you are confusing me. where is the subtraction.

jhonyy9 (jhonyy9):

= (1/3)(ln x -ln y) = = (1/3)ln x/y

jhonyy9 (jhonyy9):

bc. this is a propertie log a -log b = log a/b ok. ?

OpenStudy (baby456):

i will ask my teacher because i am confused.

OpenStudy (baby456):

ok

jhonyy9 (jhonyy9):

what is confusably for you ?

jhonyy9 (jhonyy9):

@mathmale please